A satellite is being moved from circular orbit at 400.0 km altitude to one at 250.0 km altitude. a. What is the energy of the transfer orbit? b. What is the Delta V1 needed to go from the initial circular orbit into the transfer orbit? c. What is the Delta V2 needed to go from the transfer orbit to the final circular orbit? d. What is the time (TOF) required for the transfer
Answer #1


part A :

Energy Transfer E = u/(2a)

u = GM
and

where 2a is total distance

2a = inital Height + final heght + Diamter

@a = 400+ 250 + 12742 1.34 e 7 m

so

E = -6.67 e -11* 5.9 e 24/(1.34 e 7)

E = -2.97 e 7 J


----------------------------------------------

part B dV = V2 - V1


dV = _ GMm/R1 - GMm/R2


dV = 5.97 e 8 m


----------------------------------------------------------
Time t = pi * (R1^2 + R2^2)GM

t = 0.92 secs